OpenStudy (anonymous):

$1 via paypal to the first person to answer this. [3x over y]^-3 simplify yo

OpenStudy (anonymous):

i'm not looking for money... \[(\frac{ 3x }{ y })^{-3}\] is this what you want to simplify?

OpenStudy (anonymous):

I don't want the money but since you have asked similar questions before and you still don't understand how to do this, may be you would rather prefer an explanation then a straight answer?

OpenStudy (anonymous):

yep thats what im trying to figure out!

OpenStudy (anonymous):

\[\left( a ^{d} \times b ^{e} \times c ^{f} \right) ^{g} = a ^{dg} \times b ^{eg} \times c ^{fg}\]

OpenStudy (anonymous):

Does this make sense to you?

OpenStudy (anonymous):

No, more like nauseous. Umm i tried to simplify it and got 27x^3 over y^3

OpenStudy (anonymous):

how about 27 over x^3y^3

OpenStudy (anonymous):

\[(a^b)^c = a^{b \times c}\]

OpenStudy (anonymous):

memorise this formula, you will need it.

OpenStudy (anonymous):

Now look at my previous message does it make sense to you know?

OpenStudy (anonymous):

I need a doctor, im terrible at algebra

OpenStudy (anonymous):

don't give up so easy, try to focus

OpenStudy (anonymous):

In any case, using the formula that i have written above, let's look at the problem

OpenStudy (anonymous):

(3x/y)^-3 could be written as \[\left( 3x \times \frac{ 1 }{ y } \right) ^ {-3}\]

OpenStudy (anonymous):

Now all you have to do is apply the formula to each term, first to 3x and then to 1/y

OpenStudy (anonymous):

(3x)^-3 what do you think that is?

OpenStudy (anonymous):

Lost.. what is it 3x-3 ? i got 0.9

OpenStudy (anonymous):

(3x)^-3 again apply the formula i have given you. you have 2 terms in brackets, 3 and x. Each has a power of 1 (it's just it's not written) since 3^1 = 3. You multiply -3 by 1 for each term to get 3^-3 * x^ -3 is this clear?

OpenStudy (anonymous):

I don't think i could figure this out with a pot of gold at the end of the question.

OpenStudy (anonymous):

i got 27

OpenStudy (anonymous):

for that

OpenStudy (anonymous):

What happened to x?

OpenStudy (anonymous):

well 27x then

OpenStudy (anonymous):

Let's look at this basic example more closely, I think if you can figure this (3x)^-3 out you will get the gist of the whole problem.

OpenStudy (anonymous):

27x^3?

OpenStudy (anonymous):